Rewrite as an equivalent exponential equation. Do not solve
log10^0.000001 = 6

Answer by Theo(675)

log10^0.000001 = 6

I think you might mean:

log(10,.000001) = -6 ?????

This form looks like after being put through algebra.com formula generator.

This would only be if:

10^(-6) = .000001 which looks like after being put through algebra.com formula generator.

since 10^(-6) is the same as 1/(10^6) which equals 1/1000000 which equals .000001, then I suspect this is what you are looking for.

in general, x^y = z if and only if log(x,z) = y

algebra.com formula generated form of this would look like:

in general, if and only if